# PinnDE: Physics-Informed Neural Networks for Solving Differential Equations

### v1

@@ Abstract
In recent years the study of deep learning for solving differential equations has grown substantially. The use of physics-informed neural networks (PINNs) and deep operator networks (DeepONets) have emerged as two of the most useful approaches in approximating differential equation solutions using machine learning. Here, we propose PinnDE, an open-source python library for solving differential equations with both PINNs and DeepONets. We give a brief review of both PINNs and DeepONets, introduce PinnDE along with the structure and usage of the package, and present worked examples to show PinnDE’s effectiveness in approximating solutions with both PINNs and DeepONets.
@@ 1 Introduction
Powerful numerical algorithms for solving differential equations have been studied for well over one hundred years rungekutta . Modern deep learning has been applied and studied through many diverse fields chowdhary2020natural ; voulodimos2018deep , however the use of deep learning for solving differential equations is a relatively new field. While the first ideas of these methods were proposed in lagaris1998artificial ; lagaris2000neural , in raissi2019physics the first notion of a physics-informed neural network (PINN) was popularized. Compared to finite difference methods, finite element methods or finite volume methods that use numerical differentiation for computing derivatives, using automatic differentiation baydin2018automatic in deep learning-based approaches provides a meshless alternative to solving differential equations. While raissi2019physics first proposed this for solving forward and inverse problems, many other variations of differential equations have been demonstrated to be able to be solved by PINNs. Further work on inverse problems have been disscussed in lu2021physics ; raissi2020hidden ; yu2022gradient . Solving fractional differential equations has been proposed and improved upon in the works of pang2019fpinns ; sm2024novel . Integro-differential equations have been shown to be amenable to PINN-based methods yuan2022pinn . Stochastic differential equations have been shown to be able to be solved with only slight variations to standard PINN architecture yang2020physics ; guo2022normalizing ; zhang2019quantifying . There has also been some initial work showing that including geometric properties of differential equations can improve the numerical solutions obtainable using PINNs and DeepONets, cf. arora2024invariant ; cardoso2023exactly .
@@
While there has been much research done on the different variations of differential equations solvable by PINNs, work on improving the ability of PINNs in general has been a large focus for many as well. Adaptive methods for collocation points have been a common field of study, where residual-based adaptive refinement (RAR) is a popular method which has been proposed and improved upon in the works of lu2019deeponet ; zeng2022adaptive ; nabian2021efficient ; hanna2022residual . More specific adaptive collocation point have been used in wight2020solving ; tang2022adaptive . Techniques for providing adaptive scaling parameters on the PINN architecture are developed in wang2021understanding and mcclenny2023self . Meta learning, a method for a neural network learned optimizer to optimize a PINN opposed to a hand-crafted optimizer, has been a relatively new area of research to improve PINN performance psaros2022meta ; bihlo2024improving ; liu2022novel .
@@
While PINNs are the main focus of much research around deep learning for solving differential equations, a second more general approach has grown in popularity as a tool for solving differential equations as well. Deep operator networks (DeepONets), first proposed in lu2019deeponet , provides a second general method for solving differential equation. While the basis of a PINN is to learn a particular solution of a differential equation, a DeepONet learns the solution operator for a differential equation. Thus, the network is supplied with an initial condition and appropriate boundary conditions, as well as the independent variables, rather than just the independent variables as is the case for PINNs. These networks have the ability to replace initial conditions while not having to be retrained, as well as the ability to be extended to long time intervals using time-stepping a particular solution by using the final calculated solution as a new initial condition for the learned solution operator as shown in wang2023long . Many different DeepONet variations have been proposed throughout venturi2023svd ; hadorn2022shift ; seidman2022nomad ; lee2023hyperdeeponet which show progress in the improvement in this architecture. Low order rank approximation (LoRa) hu2021lora is technique which is commonly used along with these variations to deal with the high parameter count of these networks.
@@
The Python packages TensorFlow, PyTorch, and JAX tensorflow2015-whitepaper ; pytorch ; jax2018github are commonly used for implementing deep learning algorithms and architecture. Despite this and deep learning for differential equations becoming popular research fields, there exist few software packages which implement these ideas using low-code solutions. This results in researchers and educators having to re-create already well-understood implementations from scratch. Here, we introduce PinnDE, an open-source Python library for solving differential equations with PINNs and DeepONets implemented in TensorFlow and JAX. PinnDE provides a user friendly interface in which only a minimal amount of functions is given to a user. While we find that existing packages provide powerful implementations of these algorithms, they require experience in developing software in Python, resulting in a barrier to be used by researchers or their collaborators outside of the field scientific machine learning. PinnDE provides simple to read implementations which we believe gives this package the ability to be used in education and research effectively, requiring minimal background knowledge on physics-informed machine learning by its users.
@@ 2 Background
In this section we give a short review of a physics-informed neural network (PINN) and deep operator network (DeepONet) which are the basis of the algorithms and models implemented in PinnDE.
2.1 Physics-Informed Neural Networks
@@
We consider a general initial–boundary value problem for a system of $L$ partial differential equations over a spatio-temporal domain $[t_{\rm 0},t_{\rm f}]\times\Omega$, where $\Omega\subset\mathbb{R}^{d}$, $t_{\rm 0}\in\mathbb{R}$ denotes the initial integration time, and $t_{\rm f}\in\mathbb{R}$ denotes the final integration time, given by
@@
We denote a deep neural network as $\mathcal{N}^{\theta}$ with parameters $\theta$, which includes all the weights and biases of all layers of the neural network. The goal of a physics-informed neural network is to learn to interpolate the global solution of the system of differential equations over $[t_{\rm 0},t_{\rm f}]\times\Omega$ as the parameterization $u^{\theta}=\mathcal{N}^{\theta}(t,x)$, where $u^{\theta}(t,x)\approx u(t,x)$. This is done by minimizing the loss function
$\mathcal{L}(\theta)=\mathcal{L}_{\Delta}(\theta)+\gamma_{\rm i}\mathcal{L}_{\rm i}(\theta)+\gamma_{\rm b}\mathcal{L}_{\rm b}(\theta)$ (2)
@@
This method is considered soft constrained, as the network is forced to learn the initial and boundary conditions and composite their loss, which is described in raissi2019physics . A downside of this approach is that while the initial and boundary values are known exactly, they will in general not be satisfied exactly by the learned neural network owing to being enforced only through the loss function. An alternative strategy is to hard-constrain the network with the initial and/or boundary conditions, following what was first proposed in lagaris1998artificial , which structures the neural network itself in a way so that the output of the network automatically satisfies the initial and/or boundary conditions. As such, only the differential equation loss then has to be minimized and the initial and boundary loss components in (2) are not required. The general idea of hard-constraining the initial and boundary conditions is hence to design a suitable ansatz for the neural network surrogate solution that makes sure that the initial and boundary conditions are exactly satisfied for all values of the trainable neural network. The precise form of the class of suitable hard constraints depends on the particular form of the initial and boundary value operators. We show some specific implementations in Section 3.1.1. A more formalized discussion can also be found in brecht2023improving . For many problems, it has been shown that hard constraining the initial and boundary conditions improves the training performance of neural network based differential equations solvers, and leads to lower overall errors in the obtained solutions of these solvers, cf. brecht2023improving . One notable exception are cases where the solution of a differential equation is not differentiable everywhere, in which case soft constraints typically outperform hard constraints, see e.g. brecht2024physics .
@@
We first consider a system of $L$ ordinary differential equations over the temporal domain $[t_{\rm 0},t_{\rm f}]$,
@@
where the associated loss function for the physics-informed neural network is represented as the composite loss of Eqn. (3a) and Eqn. (3b), with $\Delta^{l}$ and $\mathsf{I}^{l_{\rm i}}$ corresponding to Eqn. (4) and Eqn. (5).
One can also consider a boundary value problem with the boundary conditions
@@ 2.2 Deep Operator Networks
Deep operator networks (DeepONets) were first introduced in lu2019deeponet based on theoretical results of chen1995universal . This idea replaces learning of a particular solution of the system of differential equations with learning the solution operator acting on a specific initial–boundary condition for the system of differential equations (1). That is, let $\mathsf{G}(u_{0})(t,x)$ be a solution operator for (1) acting on the particular initial condition $u_{0}$ yielding the solution of the the initial–boundary value at the spatio-temporal point $(t,x)$, i.e. $u(t,x)=\mathsf{G}(u_{0})(t,x)$. That is, in comparison to standard physics-informed neural networks, deep operator networks require two inputs, the spatio-temporal point where the solution should be evaluated (this is the same input as for physics-informed neural networks), and the particular initial (or boundary) condition to which the solution operator should be applied. Since the initial condition is a function, it has to first be sampled on a finite dimensional subspace to be included into the neural network. The most prominent way to accomplish this is to sample a finite collection of $N_{\rm s}$ sensor points $\{x_{i}\}_{i=1}^{N_{\rm s}},x_{i}\in\Omega$ and evaluate the initial condition at those sensor points, yielding the tuple $(u_{0}(x_{1}),\dots,u_{0}(x_{N_{\rm s}}))$ that is used as an input for the deep operator network. A similar strategy is followed for pure boundary value problems. A deep operator network then uses two separate networks to combine the sampled initial conditions and independent variables. The branch network processes the initial condition sampled values, and the trunk network processes the independent variables. In the vanilla deep operator network the output vectors of the two sub-networks are then combined via a dot product. The output of deep operator network can be expressed as
@@
With $\mathcal{B}^{\theta}=(\mathcal{B}_{1}^{\theta},\dots,\mathcal{B}_{p}^{\theta})$ representing the output of the branch net and $\mathcal{T}^{\theta}=(\mathcal{T}_{1}^{\theta},\dots,\mathcal{T}_{p}^{\theta})$ representing the output of the trunk net, and $p$ being a hyper-parameter. The loss function associated with these networks is the same as (2) when soft constrained. Analogously to physics-informed neural networks, also deep operator networks can be hard-constrained with either standard PINN hard constraining methods, or further modified methods designed for deep operator networks, cf. brecht2023improving . A trained deep operator network thus is a solution interpolant for the solution operator, i.e. $\mathcal{G}^{\theta}\approx\mathsf{G}$.
@@
These networks come with multiple main benefits. Firstly, these networks can easily replace one initial condition with a new condition and do not require re-training. Secondly, as presented in wang2023long , since these networks learn the solution operator corresponding to an initial condition, they can readily be used for time-stepping and thus extend to long time intervals and maintain the ability to further predict the solution, unlike a standard PINN, which typically fails if $t_{\rm f}\gg 0$.
3 PinnDE

### v2

@@ Abstract
In recent years the study of deep learning for solving differential equations has grown substantially. The use of physics-informed neural networks (PINNs) and deep operator networks (DeepONets) have emerged as two of the most useful approaches in approximating differential equation solutions using machine learning. Here, we introduce PinnDE, an open-source Python library for solving differential equations with both PINNs and DeepONets. We give a brief review of both PINNs and DeepONets, introduce PinnDE along with the structure and usage of the package, and present worked examples to show PinnDE’s effectiveness in approximating solutions of systems of differential equations with both PINNs and DeepONets.
@@ 1 Introduction
Powerful numerical algorithms for solving differential equations have been developed for well over one hundred years rungekutta . Modern deep learning has been applied and studied through many diverse fields chowdhary2020natural ; voulodimos2018deep , however the use of deep learning for solving differential equations is a relatively new field. While the first ideas of these methods were proposed in lagaris1998artificial ; lagaris2000neural , in raissi2019physics the notion of a physics-informed neural network (PINN) was popularized. Compared to finite difference methods, finite element methods or finite volume methods that use numerical differentiation for computing derivatives, using automatic differentiation baydin2018automatic in deep learning-based approaches provides a meshless alternative to solving differential equations. While raissi2019physics first proposed this for solving forward and inverse problems, many other variations of differential equations have been demonstrated to be able to be solved by PINNs. PINNs can not only be used for solving differential equations directly, the so-called forward problem, they can also be used for parameter identification or learning of differential equations itself from data, which is referred to as the inverse problem, as discussed in lu2021physics ; raissi2020hidden ; yu2022gradient . Solving fractional differential equations has been proposed and improved upon in the works of pang2019fpinns ; sm2024novel . Integro-differential equations have been shown to be amenable to PINN-based methods yuan2022pinn as well. Stochastic differential equations were tackled with only slight variations to standard PINN architecture yang2020physics ; guo2022normalizing ; zhang2019quantifying . There has also been some initial work indicating that including geometric properties of differential equations can improve the numerical solutions obtainable using physics-informed machine learning, cf. arora2024invariant ; cardoso2023exactly .
@@
While there has been much research done on the different variations of differential equations solvable by PINNs, work on improving the ability of PINNs in general has been a large focus over the past several years as well. Adaptive methods for collocation point sampling have been a common field of study, where residual-based adaptive refinement (RAR) is a popular method which has been proposed and improved upon in the works of lu2019deeponet ; zeng2022adaptive ; nabian2021efficient ; hanna2022residual ; wu2023comprehensive . More specific adaptive collocation point have been used in wight2020solving ; tang2022adaptive . Techniques for providing adaptive scaling parameters on the PINN architecture are developed in wang2021understanding and mcclenny2023self . Meta learning, a method for a neural network learned optimizer to optimize a PINN as opposed to a hand-crafted optimizer, has been a relatively new area of research to improve physics-informed machine learning for solving differential equations psaros2022meta ; bihlo2024improving ; liu2022novel .
@@
While PINNs are the main focus using deep learning for solving differential equations, a second, more general approach has grown in popularity as a tool for solving differential equations as well. Deep operator networks (DeepONets), first proposed in lu2019deeponet , provide a general method for solving differential equations. A key distinction separates PINNs from DeepONets. While a PINN is trained to approximate the solution function, $u(x)$, for a single instance of a differential equation, a DeepONet is designed to learn the underlying solution operator, $\mathsf{G}$. The goal is to approximate the mapping from an input function $f$ (such as the initial condition, boundary conditions, or forcing term) to the corresponding solution function, such that $u(x)\approx\mathsf{G}(f)(x)$. This operator learning framework provides a significant advantage: Once trained, the DeepONet can be evaluated on new, unseen input functions without requiring retraining, making it highly efficient for solving parametric problems. Furthermore, the learned operator can be applied sequentially for long time domain predictions. By using the predicted solution at one time step as the new initial condition for the subsequent step, the network can perform time-stepping simulations, as demonstrated by wang2023long . Many different DeepONet variations have been proposed throughout venturi2023svd ; hadorn2022shift ; seidman2022nomad ; lee2023hyperdeeponet ; zeudong2025low which show progress in the improvement in this architecture.
@@
The Python machine learning frameworks TensorFlow, PyTorch, and JAX tensorflow2015-whitepaper ; pytorch ; jax2018github are commonly used for implementing deep learning algorithms and architecture. Despite this and deep learning for differential equations becoming popular research fields, there exist few software packages which implement these ideas using low-code solutions. While existing packages like DeepXDE lu2021deepxde provide powerful and flexible frameworks, they often demand significant software development experience. This results in researchers and educators having to re-create already well-understood implementations from scratch. Here, we introduce PinnDE, an open-source Python library for solving differential equations with PINNs and DeepONets implemented in TensorFlow and JAX. PinnDE provides a user-friendly interface in which only a minimal amount of functions is given to a user. PinnDE also provides simple to read implementations which we believe gives this package the ability to be used in education and research effectively, requiring minimal background knowledge on physics-informed machine learning by its users.
@@ 2 Background
In this section we give a short review of a physics-informed neural network (PINN), deep operator network (DeepONet), and inverse PINNs which are the basis of the algorithms and models implemented in PinnDE. We then give an overview of current adaptive point sampling strategies and optimization routines which PinnDE covers.
2.1 Physics-informed neural networks for forward problem
@@
We consider a general initial–boundary value problem for a system of $L$ partial differential equations over a spatio-temporal domain $t_{\rm 0},t_{\rm f}]\times\Omega$, where $\Omega\subset\mathbb{R}^{d}$, $t_{\rm 0}\in\mathbb{R}$ denotes the initial integration time, and $t_{\rm f}\in\mathbb{R}$ denotes the final integration time, given by
@@
We denote a deep neural network as $\mathcal{N}^{\theta}$ with parameters $\theta$, which includes all the weights and biases of all layers of the neural network. The goal of a physics-informed neural network is to learn to interpolate the global solution of the system of differential equations over $t_{\rm 0},t_{\rm f}]\times\Omega$ as the parameterization $u^{\theta}(t,x)=\mathcal{N}^{\theta}(t,x)$, where $u^{\theta}(t,x)\approx u(t,x)$. This is done by minimizing the loss function
$\mathcal{L}(\theta)=\mathcal{L}_{\Delta}(\theta)+\gamma_{\rm i}\mathcal{L}_{\rm i}(\theta)+\gamma_{\rm b}\mathcal{L}_{\rm b}(\theta),$ (2)
@@
This method is considered soft-constrained, as the network is forced to learn the initial and boundary conditions and composite their loss, which is described in raissi2019physics . A downside of this approach is that while the initial and boundary values are known exactly, they will in general not be satisfied exactly by the learned neural network owing to being enforced only through minimizing the loss function. An alternative strategy is to hard-constrain the network with the initial and/or boundary conditions, following what was first proposed in lagaris1998artificial , which structures the neural network itself in a way so that the output of the network automatically satisfies the initial and/or boundary conditions. As such, only the differential equation loss then has to be minimized and the initial and boundary loss components in (2) are not required. The general idea of hard-constraining the initial and boundary conditions is hence to design a suitable ansatz for the neural network surrogate solution that makes sure that the initial and boundary conditions are exactly satisfied for all values of the trainable neural network. The precise form of the class of suitable hard constraints depends on the particular form of the initial and boundary value operators. We show some specific implementations in Section 3.1.1. A more formalized discussion can also be found in brecht2023improving . For many problems, it has been shown that hard-constraining the initial and boundary conditions improves the training performance of neural network based differential equations solvers, and leads to lower overall errors in the obtained solutions of these solvers, cf. brecht2023improving . One notable exception are cases where the solution of a differential equation is not differentiable everywhere, in which case soft constraints typically outperform hard constraints, see e.g. brecht2024physics .
@@
We first consider a system of $L$ ordinary differential equations over the temporal domain $t_{\rm 0},t_{\rm f}]$,
@@
where the associated loss function for the physics-informed neural network is represented as the composite loss of Eqn. (3a) and Eqn. (3b), with $\Delta^{l}$ and $\mathsf{I}^{l_{\rm i}}$ corresponding to Eqn. (4) and Eqn. (5). One can also consider a boundary value problem with the boundary conditions
@@
2.2 Physics-informed neural networks for inverse problems
Physics-informed neural networks for inverse problems were introduced in raissi2019physics , and resemble closely the structure outlined in Section 2.1, and the description of PINN methods in the previous Section 2.1 holds true for inverse PINNs as well. The key distinction is for a deep neural network to learn to interpolate the global solution of a system of differential equations as well as parameters given in the system of differential equations. This is done by using experimental data to approximate said parameters.
In more detail, we consider the same system of partial differential equations described by Eqn. (1), with the addition of a parameter term
$\Delta^{l}(t,x,u_{(n)},\lambda)=0,\quad l=1,\dots,L,\quad t\in[t_{\rm 0},t_{\rm f}],\,x\in\Omega,$ (7a)
$\mathsf{I}^{l_{\rm i}}(x,u_{(n_{\rm i})},\lambda|_{t=t_{\rm 0}})=0,\quad l_{\rm i}=1,\dots,L_{\rm i},\quad x\in\Omega,$ (7b)
$\mathsf{B}^{l_{\rm b}}(t,x,u_{(n_{\rm b})},\lambda)=0,\quad l_{\rm b}=1,\dots,L_{\rm b},\quad t\in[t_{\rm 0},t_{\rm f}],\,x\in\partial\Omega,$ (7c)
where $\lambda=(\lambda_{1},\dots,\lambda_{p})$ denotes the tuple of parameters in this system of differential equations, with all other definitions given in Section 2.1. These parameters may also be functions of the independent variables, the dependent variables as well as the derivatives of the dependent variables.
In this case, for a neural network $\mathcal{N}^{\theta}$ with weights and biases $\theta$, the goal of an inverse PINN is to learn to interpolate the global solution of the system of differential equations over $t_{\rm 0},t_{\rm f}]\times\Omega$ and identify the values of the parameters of the equations as the parameterization $(u^{\theta},\lambda^{\theta})=(\mathcal{N}^{\theta}_{1},\dots,\mathcal{N}^{\theta}_{q+p})=\mathcal{N}^{\theta}(t,x)$, where $u^{\theta}(t,x)\approx u(t,x)$, and $\lambda^{\theta}\approx\lambda$.
This is done by adding a loss term for the experimental data to the loss function given by Eqn. (2). For a system of $L$ partial derivative equations, we add the term $\mathcal{L}_{\rm d}(\theta)$ given by
$\mathcal{L}_{\rm d}(\theta)=\frac{1}{N_{\rm d}}\sum_{i=1}^{N_{\rm d}}\sum_{k=1}^{q}|\mathcal{N}_{k}^{\theta}(t_{\rm d}^{i},x_{\rm d}^{i})-u^{k}(t_{\rm d}^{i},x_{\rm d}^{i})|^{2},$ (8)
where the network is evaluated over the collected data points sampled at the spatio-temporal data collocation points $\{(t_{\rm d}^{i},x_{\rm d}^{i})\}_{i=1}^{N_{\rm d}}$, with $N_{\rm d}$ denoting the number of data points sampled. For each of the solution components of the neural network approximation $\mathcal{N}_{k}^{\theta}(t,x)$, $k=1,\dots,q$, the mean squared error between this approximate solution and the experimental data of the underlying true solution $u^{k}(t,x)$ sampled at the data point locations is computed. In this setting, $\mathcal{N}^{\theta}$ learns to not only obtain a particular solution to a parameterized differential equation but to also fit the parameters $\lambda$ to the experimental data samples.
@@ 2.3 Deep Operator Networks
Deep operator networks (DeepONets) were first introduced in lu2019deeponet based on theoretical results of chen1995universal . This idea replaces learning of a particular solution of the system of differential equations with learning the solution operator acting on a specific initial–boundary condition for the system of differential equations (1). That is, let $\mathsf{G}(u_{0})(t,x)$ be a solution operator for (1) acting on the particular initial condition $u_{0}$ yielding the solution of the initial–boundary value at the spatio-temporal point $(t,x)$, i.e. $u(t,x)=\mathsf{G}(u_{0})(t,x)$. That is, in comparison to standard physics-informed neural networks, deep operator networks require two inputs, the spatio-temporal point where the solution should be evaluated (this is the same input as for physics-informed neural networks), and the particular initial (or boundary) condition to which the solution operator should be applied. Since the initial condition is a function, it first has to be sampled on a finite dimensional subspace to be included into the neural network. The most prominent way to accomplish this is to sample a finite collection of $N_{\rm s}$ sensor points $\{x_{i}\}_{i=1}^{N_{\rm s}},x_{i}\in\Omega$ and evaluate the initial condition at those sensor points, yielding the tuple $(u_{0}(x_{1}),\dots,u_{0}(x_{N_{\rm s}}))$ that is used as an input for the deep operator network. A similar strategy is followed for pure boundary value problems. A deep operator network then uses two separate networks to combine the sampled initial conditions and independent variables. The branch network processes the initial condition sampled values, and the trunk network processes the independent variables. In the vanilla deep operator network the output vectors of the two sub-networks are then combined via a dot product. The output of deep operator network can be expressed as
@@
With $\mathcal{B}^{\theta}=(\mathcal{B}_{1}^{\theta},\dots,\mathcal{B}_{p}^{\theta})$ representing the output of the branch net and $\mathcal{T}^{\theta}=(\mathcal{T}_{1}^{\theta},\dots,\mathcal{T}_{p}^{\theta})$ representing the output of the trunk net, and $p$ being a hyper-parameter. The loss function associated with these networks is the same as (2) when soft-constrained. Analogously to physics-informed neural networks, also deep operator networks can be hard-constrained with either standard PINN hard constraining methods, or further modified methods designed for deep operator networks, cf. brecht2023improving . A trained deep operator network thus is a solution interpolant for the solution operator, i.e. $\mathcal{G}^{\theta}\approx\mathsf{G}$.
@@
These networks come with multiple main benefits. Firstly, these networks can easily replace one initial (and/or boundary) condition with a new initial (and/or boundary) condition and do not require re-training. Secondly, as presented in wang2023long , since these networks learn the solution operator corresponding to an initial condition, they can readily be used for time-stepping for autonomous systems of differential equations and thus extend to long time intervals and maintain the ability to further predict the solution, unlike a standard PINN, which typically fails if $t_{\rm f}\gg t_{0}$. A downside of DeepONets is that they are typically harder to train as PINNs, since not only a single particular solution has to be interpolated, but the entire solution operator. This also places additional burdens on the computational hardware required, making DeepONets more expensive to train than PINNs.
2.4 Adaptive point sampling
The loss function described by Eqn. (2) relies on sampled collocation points across the domain of the system of PDEs. This is most commonly done using a random or quasi-random sample of the domain. One of the most frequently used sampling strategies is Latin hypercube sampling (LHS) mckay2000comparison which is the sampling strategy used in the first paper re-popularizing PINNs raissi2019physics .
Random sampling methods such as LHS sample collocation points independently of the solution or domain. While these methods are shown to be effective, methods of sampling collocation points which take specific solution or domain features into consideration when sampling are known as adaptive point sampling strategies. This covers a wide spectrum of strategies. Adaptive sampling strategies may include sampling strategies which add collocation points during training lu2021deepxde , resample the entire collocation point distribution during training wu2023comprehensive , or moving certain collocation points throughout the training process hou2023enhancing .
Frequently used adaptive sampling strategies in PINNs are residual-based adaptive strategies. These strategies use the PDE residual of collocation points as the factor which informs the sampling strategy. We give a brief overview of three of these strategies, where more details can be found in lu2021deepxde ; wu2023comprehensive .
Residual-based adaptive refinement (RAR) lu2021deepxde is the first of these strategies. The goal of RAR is to improve the original collocation point distribution by adding points to areas of large PDE residual. After training for $i$ epochs, a new sample of collocation points is generated. Then the PDE residual of the new points are calculated, with the $n$ points with the largest residual added to the original distribution. Here, $i,n$ are some predetermined number of iterations and collocation points, respectively.
Residual-based adaptive distribution (RAD) wu2023comprehensive is a strategy in which the entire collocation point distribution is resampled every $i$ epochs. This is done by first sampling a larger collocation point sample than the original distribution randomly. Then from this distribution we sample points to the same size of the original distribution by the probability density function
$p(x)\propto\frac{\epsilon(x)}{\mathbb{E}[\epsilon(x)]}+c$ (9)
which is computed for every collocation point $x$. Here, $\epsilon(x)$ refers to the PDE residual, where $\epsilon(x)=|\Delta(x;u_{(n)})|$. Hyperparameters $k,c$ are chosen based the desired weighting of the probability distribution function to the collocation points with higher PDE residuals.
Residual-based adaptive refinement with distribution (RAR-D) wu2023comprehensive is based on combining these two approaches. In this, after every $i$ epochs of training, we add $n$ new collocation points to the original distribution, similar to RAR. However, we add points not based on the PDE residual, but determined by the probability density function described by Eqn. (9).
3 Physics-informed neural networks for differential equations (PinnDE)